A 2.00 kg mass has forces F_1 and F_2 exerted on i
Answer #1

initiallywe need to balance forces

in x direction => F1 cos36.87 = 8N

in y direction => F2 + mg - F1 sin36.87 = 8 + 20 - 6 = 22N

therefore x component of F3 = 8N in -x direction

and y component of F3 = 22N in y direction

after halving the F3

x component of F3 = 4N in -x direction

and y component of F3 = 11N in y direction

net force in x direction = 4N in +x direction and

in -y direction net force = 11N

net force = sqrt(11^2 + 4^2) = 11.704N

accl. = 11.704/2 = 5.852 m/s2